How much energy, in the form of ATP, is obtained from b-oxidation of 1 mol of linoleic acid?


1
Expert's answer
2021-05-06T05:31:51-0400

answer:

Molecular Structure of linoleic acid:





Number of FADH2 = (n/2) - 1 - 2 double bonds

= (18/2) -1 -2

= 9 -3

= 6

"\because" 8NADH + 6FADH2 + 9Acetyl CoA

No. of ATP = 8(2.5) + 6 (1.5) + 9(10)

= (20 + 9 + 90)ATP

= 119ATP

"\therefore" Beta-oxidation of 1 mol of linoleic acid yields 119ATP